What is the minimum mass of sodium bicarbonate that would be required to completely neutralize 100.0 mL of 0.500 M sulfuric acid?
Answer #1

no of moles of H2SO4 = molarity* volume in L

                                     = 0.5*0.1 = 0.05 moles

2NaHCO3(aq) +H2SO4(aq) ---------> Na2SO4(aq) + 2CO2(g) + 2H2O(l)

1 mole of H2SO4 react with 2 moles of NaHCO3

0.05 moles of H2SO4 react with = 2*0.05/1 = 0.1 moles of NaHCO3

mass of NaHCO3 = no of moles * gram molar mass

                               = 0.1*84   = 8.4g >>>answer